OpenStudy (anonymous):

If a pebble is dropped into a pond in the shape of an ellipse at the location of one focus, the waves will converge at the location of the other focus. If the pond has a major axis of 20 feet and a minor axis of 16 feet, how far apart are the foci? 12 feet 24 feet 36 feet 6 feet

OpenStudy (anonymous):

@RadEn

OpenStudy (anonymous):

@amistre64 @mlamott15

OpenStudy (anonymous):

Sorry, I was away from my computer for a minute.

OpenStudy (anonymous):

i have no idea......what class is this from??

OpenStudy (anonymous):

Algebra 2. I'm stumped!

OpenStudy (anonymous):

@javikun1

OpenStudy (amistre64):

they are c+c apart of course ....

OpenStudy (amistre64):



OpenStudy (anonymous):

Would it be 36?

OpenStudy (amistre64):

since a is half the width; and b is half the height ... pythag for c and solve

OpenStudy (anonymous):

a = 20/2 = 10 b = 16/2 = 8 a^2 - b^2 = 10^2 - 8^2 = 100 - 64 = 36 sqrt(a^2-b^2) = sqrt(36) = 6 2*6 = X

OpenStudy (amistre64):

it has to be less than the width .... since 36 < 20 is simply absurd, then it aint 36

OpenStudy (anonymous):

6*2=X

OpenStudy (anonymous):

what is 6*2?

OpenStudy (anonymous):

I did 20^2+16^2=25.6. I don't know what I did wrong but your work makes sense to me now @Mason_Jones573

OpenStudy (amistre64):

you didnt half them ... and you added the results instead of working the pythag

OpenStudy (anonymous):

Yep^

OpenStudy (anonymous):

Oh okay, I just reworked it and got the right answer. Thank you!